# Human-in-the-Loop Generation of Adversarial Texts: A Case Study on Tibetan Script

### v3

@@ 4 Case Study
In this section, we go through the entire pipeline under the existing conditions to construct the first adversarial robustness benchmark on Tibetan script and conduct the adversarial robustness evaluation on Tibetan foundation models. We will introduce the existing conditions and the whole process in the following two subsections respectively.
@@ 4.1 Existing Conditions
Below is the involved foundation models, downstream datasets, and attack methods.
4.1.1 Foundation Models
@@ 4.1.2 Downstream Datasets
  • TNCC-title [^4] https://github.com/FudanNLP/Tibetan-Classification (Qun et al., 2017). A Tibetan news title classification dataset. It is collected from the China Tibet Online website. This dataset contains a total of 9,276 Tibetan news titles, which are divided into 12 classes.
  • TU_SA [^5] https://github.com/UTibetNLP/TU_SA (Zhu et al., 2023). A Tibetan sentiment analysis dataset. It is built by translating and proofreading 10,000 sentences from two public Chinese sentiment analysis datasets. In this dataset, negative or positive class each accounts for 50%.
@@ 4.1.3 Attack Methods
  • TSAttacker (Cao et al., 2023). An embedding-similarity-based Tibetan textual adversarial attack method. It utilizes the cosine distance between static syllable embeddings to generate substitution syllables.
  • TSTricker (Cao et al., 2024). A context-aware-based Tibetan textual adversarial attack method. It utilizes two BERT-based masked language models with tokenizers of two different granularity to generate substitution syllables or words respectively.
  • TSCheater (Cao et al., 2025). A visual-similarity-based Tibetan textual adversarial attack method. It utilizes a self-constructed Tibetan syllable visual similarity database to generate substitution candidates.
@@ 4.2 Whole Process
Figure 2 and Section 3 introduce the four stages of HITL-GAT. Below we use a case study on Tibetan script to illustrate the whole process, which is also demonstrated in Figure 3 and the video.
@@
In the victim model construction stage, we choose the foundation model and downstream dataset, and then the default fine-tuning hyperparameters will be loaded. Once the “Start” button is clicked, the fine-tuning starts and the GUI displays a progress bar, metric plots (F1/macro-F1, Accuracy, and Loss) and running logs. Here, we fine-tune Tibetan-BERT and CINO series on the training set of TNCC-title and TU_SA to construct the victim language models.
@@
Next, in the adversarial example generation stage, we choose the foundation model and downstream dataset, and then the victim language model will be loaded. Once the “Start” button is clicked, the attack starts and the GUI displays generated examples. Here, we implement TSAttacker, TSTricker, and TSCheater on the victim language models constructed from Tibetan-BERT upon the test set of TNCC-title and TU_SA to generate the first-round adversarial texts.
@@
Thereafter, in the high-quality benchmark construction stage, we screen out the examples that do not satisfy the customized filter condition $levenshtein\_distance/text\_length<=0.1$ from the first-round adversarial texts, and then manually annotate the remaining examples to construct the first Tibetan adversarial robustness benchmark called AdvTS. Given an original text and an adversarial text, we ask 3 annotators to score the human acceptance of the adversarial text based on the visual and semantic similarity between the two texts, from 1 to 5. The higher the score, the higher the human acceptance. If all annotators score the human acceptance of the adversarial text as 4 or 5, the adversarial text will be included in AdvTS.
@@
Finally, in the adversarial robustness evaluation stage, we utilize AdvTS to evaluate the adversarial robustness of CINO series with Equation 1.
@@
While a new foundation model, downstream dataset, or textual adversarial attack method emerges, we can enter the loop again to make the adversarial robustness benchmark evolve.
More case study details are given in Appendix C, including information of datasets, hyperparameters of fine-tuning, performance of victim language models, performance of textual adversarial attacks, guidelines for human annotation, etc.
5 Discussion and Limitations
The discussion and limitations are elaborated in Appendix A and B respectively.
6 Conclusion and Future Work
In this paper, we introduce a general approach and an interactive system HITL-GAT for human-in-the-loop generation of adversarial texts. Additionally, we utilize HITL-GAT to make a case study on Tibetan script. We hope that the approach and system can provide an early paradigm for constructing and updating high-quality adversarial robustness benchmarks. We also hope that the case study can serve as a reference for the adversarial research of other less-studied languages.
In the future, we will expand the functionality and improve the interaction of HITL-GAT. Also, we will use HITL-GAT to conduct more case studies on other tasks and other Chinese minority languages.
Ethics Statement
@@ 6 Conclusion and Future Work
The purpose of this paper is to promote research on the adversarial robustness of NLP models. The textual adversarial attack methods mentioned in this paper must be used positively, thus preventing any malicious misuse. Additionally, adherence to the model or dataset license is mandatory when using our system for fine-tuning, thus preventing any potential misuse.
@@ Acknowledgments
We thank the anonymous reviewers for their insightful comments, the researchers from Harbin Institute of Technology for their valuable suggestions and the annotators from Tibet University for their great efforts.
@@
Thanks to the following open-sourced projects: OpenAttack (Zeng et al., 2021), Gradio (Abid et al., 2019), LlamaFactory (Zheng et al., 2024), Transformers (Wolf et al., 2020), Datasets (Lhoest et al., 2021), etc.
@@
This work is supported by the National Social Science Foundation of China (22&ZD035), the National Natural Science Foundation of China (61972436), the Key Project of Xizang Natural Science Foundation (XZ202401ZR0040), and the MUC (Minzu University of China) Foundation (GRSCP202316, 2023QNYL22, 2024GJYY43).
Appendix A Discussion
A.1 How to define the imperceptibility of perturbations?
Due to the fact that humans perceive texts through their eyes and brains, when the perturbed text tends to the original text in visual or semantic similarity, we consider such perturbations to be imperceptible.
A.2 How to construct imperceptible perturbations?
We believe that we can start from the following three aspects.
  • Transplanting existing general methods.
    From the perspective of semantic approximation, using synonyms for substitution is a general method. Sources of synonyms can be static word embeddings (Alzantot et al., 2018), dictionaries (Ren et al., 2019), and predictions of masked language models (Garg and Ramakrishnan, 2020).
  • Using intrinsic textual features.
    Different languages have different features inherent in their texts. For example, in abugidas (Tibetan, Hindi, Bengali, etc.), many pairs of confusable letters result in visually similar syllables (Kaing et al., 2024). Figure 4 shows the visual perturbations to abugidas.
    Figure 4: Visual perturbations to abugidas.
  • Using extrinsic encoding features.
    In the process of historical development, there are many cases of “same language with different encodings”. For example, due to the technical problems in history, there are two Tibetan coded character sets in national standards of P.R.C (basic set: GB 16959-1997 and extension set: GB/T 20542-2006, GB/T 22238-2008); due to the simplification of Chinese characters, simplified and traditional Chinese exist. Figure 5 depicts the above examples.
    Encoding issues between different languages also deserve attention. For example, the Latin letter x (U+0078) and the Cyrillic letter x (U+0445) look the same; ZWNJ (zero width non-joiner, U+200C) is used extensively for certain prefixes, suffixes and compound words in Persian, but it is invisible and useless in most other languages.
    Figure 5: Encodings of Tibetan and Chinese.
Appendix B Limitations
There are several limitations in our current research. First, the fine-tuning part in victim model construction stage is far inferior to the professional fine-tuning toolkit LlamaFactory (Zheng et al., 2024). We will refer to excellent open-sourced systems to continuously expand the functionality and improve the interaction of HITL-GAT. Second, our case study now focuses on the text classification task and Tibetan script. We will use HITL-GAT to conduct more case studies on other tasks and other Chinese minority languages in the future.
Appendix C Case Study Details
C.1 Information of Datasets
Table 1 lists the detailed information of downstream datsets, including task, number of classes, average number of letters, total number of samples, etc.
C.2 Hyperparameters of Fine-tuning
Table 2 lists the default hyperparameters of downstream fine-tuning, including batch size, epochs, learning rate, etc.
C.3 Performance of Victim Language Models
Table 3 and Table 4 list the performance of victim language models on TNCC-title and TU_SA respectively.
C.4 Performance of Textual Adversarial Attacks
Table 5 lists the performance of textual adversarial attacks: TSAttacker, TSTricker, and TSCheater. “-s” and “-w” represent syllable- and word-level attack respectively. We conduct calculation of the following metrics: ADV (Accuracy Drop Value), LD (Levenshtein Distance), and CS (Cosine Similarity).
  • ADV refers to the decrease of model accuracy post-attack compared to pre-attack, as denoted below, which is usually used to evaluate the attack effectiveness. The larger the ADV, the more effective the attack.
    $ADV=Accuracy_{~{}pre}-Accuracy_{~{}post}$
  • LD between the original text and the adversarial text is the minimum number of single-letter edits (insertions, deletions, or substitutions) required to change one into the other, as denoted below, which is usually used to evaluate the visual similarity of two texts. The smaller the LD, the higher the visual similarity.
    $LD_{x,x^{\prime}}(i,j)=\begin{cases}\max(i,j)\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad if~{}min(i,j)=0,\\ LD_{x,x^{\prime}}(i-1,j-1)\quad\quad\quad\quad\quad\quad\quad if~{}x_{i}=x^{\prime}_{j},\\ 1+\min\begin{cases}LD_{x,x^{\prime}}(i-1,j)\\ LD_{x,x^{\prime}}(i,j-1)\\ LD_{x,x^{\prime}}(i-1,j-1)\end{cases}otherwise.\end{cases}$
  • CS is the cosine of the angle between two vectors, as denoted below, which is usually used to evaluate the semantic similarity of two texts. Here, the calculation is based on the word embedding space of Tibetan-BERT. The larger the CS, the higher the semantic similarity.
    $CS(x,x^{\prime})=\frac{{\textbf{x}}\cdot{\textbf{x'}}}{||\textbf{x}||~{}||\textbf{x'}||}$
Because the difference in CS is not significant, we customize the filter condition as $levenshtein\_distance/text\_length<=0.1$.
C.5 Guidelines for Human Annotation
Given an original text and an adversarial text, we ask 3 annotators to score the human acceptance of the adversarial text based on the visual and semantic similarity between the two texts, from 1 to 5. The higher the score, the higher the human acceptance. If all annotators score the human acceptance of the adversarial text as 4 or 5, the adversarial text will be included in AdvTS. Below is the guidelines for human annotation.
  • Score 1: Definite Reject.
    Humans can intuitively perceive that the perturbations significantly alter the appearance or semantics of the original text.
  • Score 2: Reject.
    Humans can intuitively perceive that the perturbations do alter the appearance or semantics of the original text.
  • Score 3: Marginal Reject or Accept.
    Humans can intuitively perceive that the perturbations alter the appearance or semantics of the original text not too much.
  • Score 4: Accept.
    After careful observation or thought for 5 seconds, humans find that perturbations only slightly alter the appearance or semantics of the original text.
  • Score 5: Definite Accept.
    After careful observation for 5 seconds, humans can not find that perturbations alter the appearance of the original text. Or, after careful thought for 5 seconds, humans find that perturbations do not alter the semantics of the original text.
C.6 Information of AdvTS
Table 6 lists the composition information of AdvTS. The subset AdvTS.TNCC-title has a total of 345 samples, and the number of samples generated by TSAttacker, TSTricker, and TSCheater is 89, 30, and 226 respectively. The average is 115, which is 12.4% of the original test set size. The subset AdvTS.TU_SA has a total of 248 samples, and the number of samples generated by TSAttacker, TSTricker, and TSCheater is 78, 19, and 151 respectively. The average is 83, which is 8.3% of the original test set size. Figure 6 shows partial samples of AdvTS. Each sample consists of 5 fields: class, original text, adversarial text, attack, and score. The construction process of the widely used adversarial robustness benchmark AdvGLUE (Wang et al., 2021a) demonstrates that most textual adversarial attack methods are prone to generating invalid or ambiguous adversarial texts, with around 90% either changing the original semantics or hindering the annotators’ unanimity. We also come to a similar conclusion. In our opinion, it is pratical and relevant to develop methods that can automatically generate high-quality adversarial texts that are visually or semantically similar to the original texts.
C.7 Adversarial Robustness of CINO
We evaluate the adversarial robustness of CINO series on AdvTS with Equation 1. The results are shown in Table 7. AdvRobust of CINO-small-v2, CINO-base-v2, and CINO-large-v2 is 0.5609, 0.5572, and 0.5726 respectively. We call on more researchers to pay attention to the model robustness of less-studied languages.
Table 1: Information of Datasets
Dataset Task #Classes #Average Letters #Total Samples #Training Samples #Validation Samples #Test Samples
TNCC-title news title classification 12 63.6196 9,276 7,422 927 927
TU_SA sentiment analysis 2 108.1897 10,000 8,000 1,000 1,000
Table 2: Hyperparameters of Fine-tuning
Model Dataset Batch Size Epochs Learning Rate Warmup Ratio Metric for Best Model
Tibetan-BERT TNCC-title & TU_SA 32 20 5e-5 0.0 Macro-F1 & F1
CINO-small-v2 TNCC-title & TU_SA 32 40 5e-5 0.1 Macro-F1 & F1
CINO-base-v2 TNCC-title & TU_SA 32 40 5e-5 0.1 Macro-F1 & F1
CINO-large-v2 TNCC-title & TU_SA 32 40 3e-5 0.1 Macro-F1 & F1
Table 3: Model Performance on TNCC-title
Model Accuracy Macro- F1 Macro- Precision Macro- Recall Weighted- F1 Weighted- Precision Weighted- Recall
Tibetan-BERT 0.6462 0.6057 0.6251 0.5956 0.6423 0.6450 0.6462
CINO-small-v2 0.7023 0.6839 0.6918 0.6819 0.7016 0.7069 0.7023
CINO-base-v2 0.6764 0.6488 0.6523 0.6556 0.6772 0.6853 0.6764
CINO-large-v2 0.7044 0.6759 0.6898 0.6672 0.7025 0.7062 0.7044
Bold and underlined values represent the best performance;
Bold values represent the second best performance.
Table 4: Model Performance on TU_SA
Model Accuracy F1 Precision Recall
Tibetan-BERT 0.7070 0.6913 0.7305 0.6560
CINO-small-v2 0.7550 0.7818 0.7047 0.8780
CINO-base-v2 0.7530 0.7748 0.7119 0.8500
CINO-large-v2 0.7970 0.7992 0.7906 0.8080
Bold and underlined values represent the best performance;
Bold values represent the second best performance.
Table 5: Performance of Textual Adversarial Attacks
Metric Method TNCC-title TU_SA
Tibetan-BERT CINO-small CINO-base CINO-large Tibetan-BERT CINO-small CINO-base CINO-large
ADV $\uparrow$ TSAttacker 0.3420 0.3592 0.3646 0.3430 0.1570 0.2260 0.2240 0.2660
TSTricker-s 0.5124 0.5685 0.5414 0.5426 0.3080 0.4300 0.4730 0.5060
TSTricker-w 0.5124 0.5588 0.5566 0.5286 0.2870 0.4050 0.4200 0.5100
TSCheater-s 0.4714 0.5717 0.5620 0.5329 0.2810 0.3790 0.4390 0.4280
TSCheater-w 0.5027 0.5696 0.5696 0.5405 0.2810 0.3770 0.4540 0.4260
LD $\downarrow$ TSAttacker 5.2000 5.6210 5.0638 5.3386 7.7298 7.4533 8.0769 7.3369
TSTricker-s 4.0671 5.8856 5.3402 5.6865 5.4887 6.2495 9.2057 6.8813
TSTricker-w 10.2492 13.0297 12.9511 12.3374 16.9542 14.2365 16.5066 16.7699
TSCheater-s 1.6941 2.5775 2.5713 2.8846 2.1219 3.5827 4.7923 3.5393
TSCheater-w 3.1771 3.9363 4.0066 4.0531 7.4147 9.0522 10.0106 8.9047
CS $\uparrow$ TSAttacker 0.9653 0.9644 0.9678 0.9666 0.9844 0.9862 0.9841 0.9845
TSTricker-s 0.9602 0.9543 0.9603 0.9578 0.9750 0.9793 0.9739 0.9778
TSTricker-w 0.8865 0.8870 0.8895 0.8925 0.9315 0.9384 0.9316 0.9371
TSCheater-s 0.9547 0.9734 0.9737 0.9708 0.9785 0.9903 0.9865 0.9908
TSCheater-w 0.9447 0.9433 0.9433 0.9547 0.9417 0.9507 0.9501 0.9526
Bold and underlined values represent the best performance;
Bold values represent the second best performance.
Table 6: Composition Information of AdvTS
Subset TSAttacker TSTricker TSCheater Average Total
AdvTS.TNCC-title 89 30 226 115 345
AdvTS.TU_SA 78 19 151 83 248
Figure 6: Partial Samples of AdvTS
Table 7: Adversarial Robustness of CINO on AdvTS
Model $Accuracy$ on AdvTS.TNCC-title $Accuracy$ on AdvTS.TU_SA $AdvRobust$ on AdvTS
CINO-small-v2 0.4928 0.6290 0.5609
CINO-base-v2 0.4812 0.6331 0.5572
CINO-large-v2 0.5362 0.6089 0.5726

### v5

@@ 4 Case Study
In this section, we go through the entire pipeline under the existing conditions to construct the first adversarial robustness benchmark for Tibetan script and conduct the adversarial robustness evaluation on Tibetan language models. We will introduce the existing conditions and the whole process in the following two subsections respectively.
@@ 4.1 Existing Conditions
Below is the involved language models, downstream datasets, and attack methods.
4.1.1 Language Models
@@ 4.1.2 Downstream Datasets
  • TNCC-title [^4] https://github.com/FudanNLP/Tibetan-Classification (DBLP:conf/cncl/QunLQH17). A Tibetan news title classification dataset. This dataset contains a total of 9,276 Tibetan news titles, which are divided into 12 classes.
  • TU_SA [^5] https://github.com/UTibetNLP/TU_SA (MESS202302007). A Tibetan sentiment analysis dataset. It is built by translating and proofreading 10,000 sentences from two public Chinese sentiment analysis datasets. In this dataset, negative or positive class each accounts for 50%.
@@ 4.1.3 Attack Methods
Over the past few years, we have developed several Tibetan textual adversarial attack methods, aiming to draw attention to the NLP security in lower-resourced languages, as listed below. Our past work (cao-etal-2023-pay-attention) is the only one engaged with a truly low-resource language among the research samples in the literature NLP Security and Ethics, in the Wild (lent-etal-2025-nlp, TACL 2025, page 719).
  • TSAttacker (cao-etal-2023-pay-attention). An embedding-similarity-based Tibetan textual adversarial attack. It utilizes the cosine distance between static syllable embeddings to generate substitution syllables.
  • TSTricker (10.1145/3589335.3652503). A context-aware-based Tibetan textual adversarial attack. It utilizes two BERT-based masked language models with tokenizers of two different granularities to generate substitution syllables or words respectively.
  • TSCheater (10889732). A visual-similarity-based Tibetan textual adversarial attack. It utilizes a self-constructed Tibetan syllable visual similarity database to generate substitution candidates.
@@ 4.2 Whole Process
Figure 2 and Section 3 introduce the four stages of HITL-GAT. Below we use a case study on Tibetan script to illustrate the whole process, which is also demonstrated in the video and Figure 3.
Figure 3: Screenshots of HITL-GAT.
@@
In the victim model construction stage, we choose the language model and downstream dataset, and then the default fine-tuning hyperparameters will be loaded. Once the “Start” button is clicked, the fine-tuning starts and the GUI displays a progress bar, metric plots (F1/macro-F1, Accuracy, and Loss) and running logs. Here, we fine-tune Tibetan-BERT and CINO series on the training set of TNCC-title and TU_SA to construct the victim language models.
@@
Next, in the adversarial example generation stage, we choose the language model and downstream dataset, and then the victim language model will be loaded. Once the “Start” button is clicked, the attack starts and the GUI displays generated examples. Here, we implement TSAttacker, TSTricker, and TSCheater on the victim language models constructed from Tibetan-BERT upon the test set of TNCC-title and TU_SA to generate the first-round adversarial texts.
@@
Thereafter, in the high-quality benchmark construction stage, we screen out the examples that do not satisfy the customized filter condition levenshtein_distance/text_length<=0.1levenshtein\_distance/text\_length<=0.1 from the first-round adversarial texts, and then manually annotate the remaining examples to construct the first Tibetan adversarial robustness benchmark called AdvTS. Given an original text and an adversarial text, we ask 3 annotators to score the human acceptance of the adversarial text based on the visual and semantic similarity between the two texts, from 1 to 5. The higher the score, the higher the human acceptance. If all annotators score the human acceptance of the adversarial text as 4 or 5, the adversarial text will be included in AdvTS. Below is the guidelines for human annotation.
  • Score 1: Definite Reject. Humans can intuitively perceive that the perturbations significantly alter the appearance or semantics of the original text.
  • Score 2: Reject. Humans can intuitively perceive that the perturbations do alter the appearance or semantics of the original text.
  • Score 3: Marginal Reject or Accept. Humans can intuitively perceive that the perturbations alter the appearance or semantics of the original text not too much.
  • Score 4: Accept. After careful observation or thought for 5 seconds, humans find that perturbations only slightly alter the appearance or semantics of the original text.
  • Score 5: Definite Accept. After careful observation for 5 seconds, humans can not find that perturbations alter the appearance of the original text. Or, after careful thought for 5 seconds, humans find that perturbations do not alter the semantics of the original text.
@@
Finally, in the adversarial robustness evaluation stage, we utilize AdvTS to evaluate the adversarial robustness of CINO series with Equation 1. The $AdvRobust$ of CINO-small-v2, CINO-base-v2, and CINO-large-v2 is 0.5609, 0.5572, and 0.5726 respectively.
@@
While a new language model, downstream dataset, or textual adversarial attack method emerges, we can enter the loop again to make the adversarial robustness benchmark evolve.
5 Discussion
Due to the fact that humans perceive texts through their eyes and brains, when the perturbed text tends to the original text in visual or semantic similarity, we consider such perturbations to be imperceptible. To construct imperceptible perturbations, we can start from the following three aspects.
  • Transplanting existing general methods. From the perspective of semantic approximation, using synonyms for substitution is a general method. Sources of synonyms can be static word embeddings (alzantot-etal-2018-generating), dictionaries (ren-etal-2019-generating), and predictions of masked language models (li-etal-2020-bert-attack).
  • Using intrinsic textual features. Different languages have different features inherent in their texts. For example, in abugidas (Tibetan script, Devanagari script, etc.), many pairs of confusable letters result in visually similar syllables (kaing-etal-2024-robust; 10889732).
  • Using extrinsic encoding features. In the process of historical development, there are many cases of “same language with different encodings”. For example, due to the technical problems in history, there are two Tibetan coded character sets in national standards of P.R.C (basic set: GB 16959-1997 and extension set: GB/T 20542-2006, GB/T 22238-2008); due to the simplification of Chinese characters, simplified and traditional Chinese exist. Encoding issues between different languages also deserve attention. For example, the Latin letter x (U+0078) and the Cyrillic letter x (U+0445) look the same; ZWNJ (zero width non-joiner, U+200C) is used extensively for certain prefixes, suffixes and compound words in Persian, but it is invisible and useless in most other languages.
6 Conclusion
This paper introduces HITL-GAT, an interactive system for human-in-the-loop generation of adversarial texts. Our approach employs a four-stage iterative loop: victim model construction, adversarial example generation, high-quality benchmark construction, and adversarial robustness evaluation. The loop ensures adversarial robustness benchmarks to co-evolve with advancements in language models, downstream datasets, and textual adversarial attack methods. Additionally, we demonstrate the utility of HITL-GAT through a case study on Tibetan script, employing three customized adversarial text generation methods and establishing its first adversarial robustness benchmark. Our work provides a valuable reference for other lower-resourced languages, especially languages in the Asia-Pacific that use abugidas as their writing system. The weaponization of lower-resourced languages against NLP security highlights the critical gap and the urgent need for research in this area.
Limitations
The system and the case study presented in this paper are the crystallization of our research on the adversarial robustness of Tibetan language models over the past few years. The summarized approach is only applicable to classification tasks. Given the heightened sensitivity necessary for working with lower-resourced languages, our case study is conducted on insensitive tasks with ethical best practices in mind. Due to the existing conditions of lower-resourced languages, our case study can only be conducted this far. However, this does not prevent it from serving as an early paradigm for researching the evolution of adversarial robustness benchmarks. We will continue to develop HITL-GAT and conduct more case studies on other minority languages.
Ethical Considerations
@@ 6 Conclusion
Our work adheres to the ACM Code of Ethics. The purpose of this paper is to promote research on NLP security, especially for lower-resourced languages. The textual adversarial attack methods mentioned in this paper must be used positively, thus preventing any malicious misuse. Additionally, adherence to the model or dataset license is mandatory when using our system or fork versions, thus preventing any potential misuse.
@@ Acknowledgments
This work benefited greatly from the encouragement, advice, and help of others. Thank you to Heather Lent at Aalborg University for her encouragement and detailed feedback on our manuscript. Thank you to Chen Zhang at Peking University and Andong Chen at Harbin Institute of Technology for their early advice. We also thank the anonymous reviewers of *ACL for their insightful comments.
@@
Finally, thanks to the following open-sourced projects: OpenAttack (zeng-etal-2021-openattack), Gradio (abid2019gradiohasslefreesharingtesting), LlamaFactory (zheng-etal-2024-llamafactory), Transformers (wolf-etal-2020-transformers), Datasets (lhoest-etal-2021-datasets), etc.
@@
This work is supported by the Science and Technology Strategic Consulting Project of the Chinese Academy of Engineering (2025-XZ-16-06), the Key Project of Xizang Natural Science Foundation (XZ202401ZR0040), the National Social Science Foundation of China (22&ZD035), the National Natural Science Foundation of China (61972436), and the MUC (Minzu University of China) Foundation (2025XYCM39).